# GraphAr: An Efficient Storage Scheme for Graph Data in Data Lakes

### 18 Dec 2023 v2

@@ 6.1 Experimental Setup
Baselines. GraphAr is developed in C++ on Apache Arrow [2], an open-source, high-performance library that supports columnar formats like Apache Parquet [5] and Apache ORC [4]. We benchmark GraphAr against Apache Arrow/Parquet (version 13.0.0), due to the popularity and high-performance of Parquet. While there are other related works in this area, we discuss them in Section 7 and do not directly compare them with GraphAr, since they are either not applicable or demonstrate worse performance than Parquet in the context of data lakes. Both GraphAr and the baseline follow Parquet’s default configurations, such as a row group length of $1024\times 1024$ rows and a 1MB page size. Dictionary encoding is avoided due to its decoding overhead, as discussed in Section 6.3.1.
Datasets. As summarized in Table 1, our evaluation includes a variety of graphs, span different sizes and domains, including social networks and web graphs. We also use synthetic graphs generated by the LDBC SNB data generator [32], designed to mimic real-world graph characteristics. These synthetic graphs are part of the LDBC Social Network Benchmark [32, 57], and we use its query set for end-to-end workload assessments.
@@ 6.2.1 Storage Efficiency
We compare GraphAr with baseline methods by measuring the storage consumed by encoded Parquet files that store the graph’s topological data. Two baseline approaches are considered: 1) “plain”, which employs plain encoding for the source and destination columns, and 2) “plain + offset”, which extends the “plain” method by sorting edges and adding an offset column to mark each vertex’s starting edge position. As Figure 8 depicts, the inclusion of offsets results in a modest increase in storage requirements, with an increase in space usage ranging from $0.5\%$ to $14.8\%$, as the number of vertices is typically much smaller than the number of edges.
@@
Figure 8: Comparison of storage consumption for topology.
@@ 6.2.2 Performance of Neighbor Retrieval
To evaluate GraphAr’s efficiency in neighbor retrieval, we query vertices with the largest degree in selected graphs, maintaining edges in CSR-like or CSC-like formats depending on the degree type. Figure 9 shows that GraphAr significantly outperforms the baselines, achieving speedups ranging from 49.8×49.8$\times$49.8 × to 9232.4×9232.4$\times$9232.4 × over the “plain” method and 2.1×2.1$\times$2.1 × to 6.1×6.1$\times$6.1 × over the “plain + offset” baseline. These gains are attributed to the offset integration and delta encoding, as well as our proposed BMI-based decoding strategy. The offset integration alone accounts for a speedup of between 11.0×11.0$\times$11.0 × and 4177.2×4177.2$\times$4177.2 ×, and delta encoding contributes an additional 1.8×1.8$\times$1.8 × to 3.6×3.6$\times$3.6 × speedup on top of that. While not explicitly shown in the log-scaled Figure 9, the use of BMI and SIMD further enhances performance, with an average improvement of $20.1\%$.
Figure 9: Comparison of neighbor retrieval time.
@@ 6.2.3 Performance of Data Transformation
Given that GraphAr is designed for storing graph data in data lakes, the efficiency of data transformation—converting original graph data into the GraphAr format is crucial. We focus on the time required for this transformation, using four large graphs (U2, AR, U5, and WB) represented as Arrow Tables, a standardized in-memory format in big data systems. Given that graphs generally have significantly more edges than vertices and that GraphAr employs CSR/CSC-like layouts requiring edge sorting, generating topological data becomes the most time-intensive part of the process.
Figure 10: Time breakdown of data transformation.
To dissect the time overhead, we perform a breakdown analysis as shown in Figure 10. The process involves 3 key steps: 1) sorting the edges using Arrow/Acero’s internal order_by operator, labeled as “sort”; 2) generating vertex offsets, labeled as “offset”; and 3) writing the sorted and offset data into Parquet files with specific encoding, labeled as “output”. Generally, the sorting and encoding process has a time complexity of $O(|E|log|E|)$ in sequential. The WB graph takes less time than U5 because WB’s original data has clustered edges by source but not sorted by destination. On the other hand, the original data in U5 is randomly distributed.
Our experiments show that we can generate topological data for more than $3.9$ million edges per second. Considering this transformation is a one-time, offline operation that substantially reduces future data retrieval times, the associated overhead is acceptable. The most time-consuming step, edge sorting, offers room for further optimization through distributed processing frameworks like Apache Spark.
@@ 6.3 Micro-Benchmark of Label Filtering
This section evaluates GraphAr’s efficiency in storing and filtering vertex labels. We employ datasets from OGB [39] and Neo4j examples [14], which feature property graphs with multiple vertex labels ranging from 3 to 349.
@@ 6.3.1 Storage Efficiency
As shown in Figure 11, our RLE-based method substantially outperforms the baselines, requiring on average only 2.7% and 10.8% of the storage space compared to the “string” and “binary (plain)” methods, respectively. While Parquet does support dictionary encoding that could potentially enhance the “string” baseline, we excluded this from our evaluation. The reason being that, despite some storage gains, dictionary encoding incurs a decoding slowdown of up to 10x due to the extra overhead of storing the dictionary, especially when label numbers are high.
Our RLE-based strategy strikes a balance between storage efficiency and decoding performance, as demonstrated in the subsequent experiments.
Figure 11: Comparison of storage consumption for labels (the total number of labels is shown under the graph name).
@@ 6.3.2 Performance of Simple Condition Filtering
Figure 12 illustrates the results, demonstrating that the GraphAr method significantly improves the performance of label filtering based on simple conditions. The most straightforward approach, “string”, which involves decoding the string of labels and conducting matching for each vertex, is the slowest. The “binary (plain)” method separates labels into individual columns and utilizes a binary representation, while the ‘binary (RLE)” method further optimizes the encoding by using RLE. However, both of these methods still require evaluating each vertex. In contrast, our method, referred to as “binary (RLE) + interval”, simply selects all satisfied intervals.
In Figure 12, for each graph, we report the middle value of the execution time among filtering each label as the height of the bar, with the error bar representing the range of execution time. Our method may have a large range on some graphs (TV, AX, MA) due to the varying encoding efficiency (i.e., the number of intervals generated) for different columns. However, since the number of intervals is not larger than the number of vertices in any case, our method consistently outperforms the baselines. On average, it achieves a speedup of 6.0×6.0$\times$6.0 × over the “string” method, $\times3$ over the “binary (plain)” method, and 2.3×2.3$\times$2.3 × over the “binary (RLE)” method.
Figure 12: The execution time of simple condition filtering.
Figure 13: Comparison of speedup over the “string” baseline on complex condition filtering.
@@ 6.3.3 Performance of Complex Condition Filtering
In this micro-benchmark, we assess the performance of label filtering based on complex conditions using the same graphs as mentioned above. For graphs obtained from Neo4j examples, we first refer to the provided documentation to identify a filtering operation that involves two labels. If not provided, we create a condition by combining two related labels using either the logical AND operator (if there are vertices satisfying the condition) or OR (otherwise) to reflect real-world semantics. For graphs obtained from the OGB library, we utilize the first two labels and combine them using OR as the filtering condition. We then conduct experiments on the encoded Parquet files and measure the execution time to obtain the vertices that satisfy the condition.
Figure 13 presents the speedup of different methods against the most straightforward approach (referred to as “string” in Figure 11 and Figure 12). The results demonstrate that GraphAr performs the best for all test cases. Further analysis reveals that the performance improvement is attributed to the binary representation (as observed in the comparison between “binary (plain)” and “string”), and utilization of RLE (as seen in the comparison between “binary (RLE)” and “binary (plain)”). However, the most significant improvement comes from the interval-based evaluation, as observed in the comparison between “binary (RLE) + interval” and “binary (RLE)”, which results in a further speedup of up to 56.3×56.3$\times$56.3 ×.
@@ 6.3.4 Scale Up the Number of Labels
Although label filter conditions in real-world workloads usually involve only a few labels, we also evaluate the performance as the number of labels increases. Figure 14 illustrates the execution time of filtering conditions with varying numbers of labels, focusing on BL and MA. These graphs are selected from different datasets (Neo4j and OGB) and have a relatively large number of labels. To test the filtering with $i$ labels, we use the first $i$ labels in a graph and combine them through OR as the condition (to ensure that there are vertices satisfying the condition). To reduce experimental noise, we conduct the experiments for $100$ times and report the total execution time for BL and the average time for MA.
As shown in the figure, GraphAr consistently outperforms others on the BL graph. While on the MA graph, it performs best when the number of labels involved is not larger than $40$. When the number of labels continuously increases, it performs worse than the baseline “binary (RLE)”. This is because as the number of labels increases, the number of merged intervals also increases, resulting in a larger number of times the UDF is called. In the worst case, the UDF is called for each vertex, means any two consective vertices have different labels. Considering the additional overhead of merging intervals from different columns, our method may perform worse than directly evaluating the UDF for each vertex. Fortunately, our investigation of real-world workloads reveals that the number of filtered labels in user-written queries is often limited. For instance, in the documentation of Neo4j examples  [14], the filtering involves no more than $5$ labels. Consequently, the performance of our method remains highly promising.
(a) BL
(b) MA
Figure 14: Execution time with the number of labels increases.
@@ 6.4 Storage Media
We assess the efficiency of GraphAr across various storage media: local in-memory tmpfs, ESSD (an Alibaba Cloud virtualized elastic block device), and S3-like Object Store Service (OSS). The dataset used is SF100 and the comment vertex type and comment_hasTag_tag edge type are used.
@@
Table 2 encapsulates the efficiency of GraphAr across different storage media and operations. These results conclusively establish that GraphAr is not just efficient but also robust, delivering consistently high performance across storage media, with speedups of $\times8$ to 154×154$\times$154 × for neighbor retrieval and 2.7×2.7$\times$2.7 × to $\times1$ for label filtering.
6.5 End-to-end Workloads
To demonstrate the practicality of GraphAr in real-world scenarios, we conduct a performance evaluation using end-to-end workloads from the LDBC SNB benchmark [32, 57], which is widely used in the graph processing community. Upon investigating this benchmark, including 7 short and 14 complex interactive queries, as well as 20 business intelligence queries, we find that neighbor retrieval is frequently encountered, involved in approximately $90\%$ of the queries.
Although the benchmark specifies vertex/edge types, it does not explicitly define the labels. However, we are able to identify certain vertex types that are static (e.g., tagclass and place), which have a fixed and very small vertex set size that does not scale with the graph size. On the other hand, vertex types like comment and person are considered dynamic. Based on this observation, we can treat information related to static types as labels for dynamic types in GraphAr, for example, all tag classes of a comment are attached as labels for the corresponding comment vertex. Similar strategies are also adopted by graph databases [37] to optimize data access performance. Considering this organization, approximately $50\%$ of the queries involve label filtering.
For evaluation, graphs at different scales (listed in Table 1) are generated using the LDBC SNB data generator, as Parquet files. These graphs are then converted into GraphAr format, with the vertex labels attached as described above.
@@ 6.5.1 Query Implementations
Queries. The evaluation focuses on three representative queries, with the required parameters set according to the reference implementations [12, 11]. The first query, IS-3 (interactive-short-3), aims to find all the friends of a given person and return their information. It exemplifies the common pattern of querying neighboring vertices and retrieving associated properties. The second query, IC-8 (interactive-complex-8), is more complex as it involves traversing multiple hops from the starting vertex. Lastly, the BI-2 (business-intelligence-2) query aims to analyze tag evolution. It involves finding and counting the messages associated with tags within a specific tag class, thus requiring vertex filtering by labels. These selected queries represent common patterns in graph queries, and our approach is applicable to other LDBC queries, as approximately $90\%$ of the queries involve neighbor retrieval and $50\%$ involve label filtering.
GraphAr. We develop hand-written implementations for each query using the GraphAr APIs, which utilize the data organization and specifically prioritize two essential operations: neighbor retrieval and label filtering. Our implementation refers to the official reference implementations [12, 11] in Cypher or SQL, to ensure equivalence to the original queries.
Acero. We also implement the same queries using the APIs provided by Acero [1], which is a powerful C++ library integrated into Apache Arrow, designed for analyzing large streams of data. It offers a comprehensive set of operators such as scan, filter, project, aggregate, and join, among others. Moreover, Acero supports taking Parquet files as the data source and enables the pushdown of predicates. These capabilities make it an appropriate choice for serving as the baseline for comparing with GraphAr. Despite our best efforts to optimize it, we do not perform any data re-organization or utilize the encoding/decoding optimizations of GraphAr, for the implementations based on Parquet and Acero.
@@ 6.5.2 Performance Comparison
Table 3 presents a comparison of end-to-end time performance between implementations based on GraphAr and Acero. The results clearly demonstrate that GraphAr significantly outperforms Acero, achieving an average speedup of 29.5×29.5$\times$29.5 ×. Upon closer examination of the results, it becomes apparent that the performance improvement can be attributed to the following factors: 1) data layout design and encoding/decoding optimizations we proposed, to enable efficient neighbor retrieval (IS-3, IC-8, BI-2) and label filtering (BI-2), as demonstrated in micro-benchmarks; 2) bitmap generation during the operations of neighbor retrieval and label filtering, which can be utilized in subsequent selection steps (IS-3, IC-8, BI-2).
Taking IS-3 as an example, the optimization of neighbor retrieval achieves an average speedup of 571.4×571.4$\times$571.4 × compared to the baseline utilizing Parquet’s filter pushdown. The subsequent selection based on bitmaps provides an average speedup of 2.7×2.7$\times$2.7 ×. Other steps, such as order_by, exhibit nearly indistinguishable performance due to the closely matched implementations in GraphAr and Acero. Overall, these results demonstrate the practicality and effectiveness of GraphAr in end-to-end workloads.
Table 3: Execution time for end-to-end workloads IS-3, IC-8 and BI-2 (in seconds).
@@
Summary. In the comprehensive evaluation, GraphAr has been demonstrated to be a highly effective storage scheme for Labeled Property Graphs (LPGs) in data lakes. The key takeaways are:
  • Storage efficiency: GraphAr remarkably reduces storage requirements, needing on average only $27.3\%$ of the storage used by baseline methods in neighbor retrieval and as low as $2.7\%$ for label filtering.
  • Query performance: GraphAr significantly outpaces the baselines in retrieval time, achieving speedups ranging from 49.8×49.8$\times$49.8 × to 9232.4×9232.4$\times$9232.4 × for neighbor retrieval and an average speedup of 6.0×6.0$\times$6.0 × for simple label filtering, based on micro-benchmarks on the ESSD storage.
@@
  • Real-world relevance: In end-to-end workloads using the LDBC SNB benchmark, GraphAr shows an average speedup of 29.5×29.5$\times$29.5 × over the Acero baseline, substantiating its practical utility.

### 21 Dec 2023 v3

@@ 6.1 Experimental Setup
Baselines. GraphAr is developed in C++ on Apache Arrow [2], an open-source, high-performance library that supports columnar formats like Parquet and ORC. We benchmark GraphAr against Apache Arrow/Parquet (version 13.0.0), due to the popularity and high-performance of Parquet. While there are other related works in this area, we discuss them in Section 7 and do not directly compare them with GraphAr, since they are either not applicable or demonstrate worse performance than Parquet in the context of data lakes. Unless otherwise specified, both GraphAr and the baseline follow Parquet’s default configurations, which include a row group length of $1024\times 1024$ rows and a 1MB page size.
Datasets. As summarized in Table 1, our evaluation includes a variety of graphs, span different sizes and domains, including social networks and web graphs. We also use synthetic graphs generated by the LDBC SNB data generator [32], designed to mimic real-world graph characteristics. They are part of the LDBC Social Network Benchmark [32, 57], and we use its query set for end-to-end workload assessments.
@@ 6.2.1 Storage Efficiency
We compare GraphAr with baseline methods by measuring the storage consumed by encoded Parquet files that store the graph’s topological data. Two baseline approaches are considered: 1) “plain”, which employs plain encoding for the source and destination columns, and 2) “plain + offset”, which extends the “plain” method by sorting edges and adding an offset column to mark each vertex’s starting edge position. As Figure 7(a) depicts, the inclusion of offsets results in a modest increase in storage requirements, with an increase in space usage ranging from $0.5\%$ to $14.8\%$, as the number of vertices is typically much smaller than the number of edges.
@@
(a) Storage consumption for topology.
(b) Neighbor retrieval time.
(c) Data transformation time.
Figure 8: Micro-benchmark of neighbor retrieval.
@@ 6.2.2 Performance of Neighbor Retrieval
To evaluate GraphAr’s efficiency in neighbor retrieval, we query vertices with the largest degree in selected graphs, maintaining edges in CSR-like or CSC-like formats depending on the degree type. Figure 7(b) shows that GraphAr significantly outperforms the baselines, achieving speedups ranging from 49.8×49.8$\times$49.8 × to 9232.4×9232.4$\times$9232.4 × over the “plain” method and 2.1×2.1$\times$2.1 × to 6.1×6.1$\times$6.1 × over the “plain + offset” baseline. These gains are attributed to the offset integration and delta encoding, as well as our proposed BMI-based decoding strategy. The offset integration alone accounts for a speedup of between 11.0×11.0$\times$11.0 × and 4177.2×4177.2$\times$4177.2 ×, and delta encoding contributes an additional 1.8×1.8$\times$1.8 × to 3.6×3.6$\times$3.6 × speedup on top of that. While not explicitly shown in the log-scaled Figure 7(b), the use of BMI and SIMD further enhances performance, with an average improvement of $20.1\%$.
@@ 6.2.3 Performance of Data Transformation
Given that GraphAr is designed for storing LPGs in data lakes, the efficiency of converting original graph data into the GraphAr format is crucial. We focus on the time required for this transformation, using four large graphs (U2, AR, WB and U5) represented as Arrow Tables, a standardized in-memory format in big data systems. Since graphs generally have significantly more edges than vertices and GraphAr employs CSR/CSC-like layouts requiring edge sorting, generating topological data becomes the most time-intensive part.
To dissect this time overhead, we perform a breakdown analysis in Figure 7(c), which shows that we can generate topological data for over $3.9$ million edges per second. The process involves three steps: 1) sorting the edges using Arrow/Acero’s internal order_by operator, labeled as “sort”; 2) generating vertex offsets, labeled as “offset”; and 3) writing the sorted and offset data into Parquet files with specific encoding, labeled as “output”. Generally, this process has a time complexity of $O(|E|log|E|)$ sequentially. Considering this transformation is a one-time, offline operation that substantially reduces future data retrieval times, the associated overhead is acceptable. The most time-consuming step, sorting, offers room for further optimization through distributed frameworks like Spark.
@@ 6.3 Micro-Benchmark of Label Filtering
This section evaluates GraphAr’s efficiency in storing and filtering vertex labels. We employ datasets from OGB [39] and Neo4j [14], which feature property graphs with multiple vertex labels, with the number of labels (ranging from $3$ to $349$) indicated under the graph name in Figure 8(a).
@@ 6.3.1 Storage Efficiency
As shown in Figure 8(a), our RLE-based method substantially outperforms the baselines, requiring on average only $2.9\%$ and $10.1\%$ of the storage space compared to the “string” and “binary (plain)” methods, respectively. While Parquet does support dictionary encoding that could potentially enhance the “string” baseline, we excluded this from our evaluation. The reason being that, despite some storage gains, dictionary encoding incurs a decoding slowdown of up to 10×10$\times$10 × due to the extra overhead of storing the dictionary, especially when label numbers are high. Our RLE-based strategy strikes a balance between storage efficiency and decoding performance, as demonstrated in the subsequent experiments.
@@ 6.3.2 Performance of Simple Condition Filtering
Figure 8(b) illustrates the results, demonstrating that the GraphAr method significantly improves the performance of label filtering based on simple conditions. The most straightforward approach, “string”, which involves decoding the string of labels and conducting matching for each vertex, is the slowest. The “binary (plain)” method separates labels into individual columns and utilizes a binary representation, while the ‘binary (RLE)” method further optimizes the encoding by using RLE. However, both of these methods still require evaluating each vertex. In contrast, our method, referred to as “binary (RLE) + interval”, simply selects all satisfied intervals.
In Figure 8(b), for each graph, we report the middle value of the execution time among filtering each label as the height of the bar, with the error bar representing the range of execution time. Our method may have a large range on some graphs (AX, MA) due to the varying encoding efficiency (i.e., the number of intervals generated) for different columns. However, since the number of intervals is not larger than the number of vertices in any case, our method consistently outperforms the baselines. On average, it achieves a speedup of 6.0×6.0$\times$6.0 × over the “string” method, $\times3$ over the “binary (plain)” method, and 2.3×2.3$\times$2.3 × over the “binary (RLE)” method.
(a) Storage consumption for labels.
(b) Simple condition filtering.
(c) Complex condition filtering.
Figure 9: Micro-benchmark of label filtering.
@@ 6.3.3 Performance of Complex Condition Filtering
We also assess the performance of label filtering based on complex conditions using the same graphs as mentioned above. For graphs obtained from Neo4j, we first refer to the provided documentation to identify a filtering operation that involves two labels. If not provided, we create a condition by combining two related labels using either the logical AND operator (if there are vertices satisfying the condition) or OR (otherwise) to reflect real-world semantics. For graphs obtained from OGB, we utilize the first two labels and combine them using OR as the filtering condition.
Figure 8(c) presents the speedup of different methods against the most straightforward approach (referred to as “string” in Figure 8(a) and Figure 8(b)). The results demonstrate that GraphAr performs the best for all test cases. Further analysis reveals that the performance improvement is attributed to the binary representation (as observed in the comparison between “binary (plain)” and “string”), and utilization of RLE (as seen in the comparison between “binary (RLE)” and “binary (plain)”). However, the most significant improvement comes from the merge-based decoding, as observed in the comparison between “binary (RLE) + merge” and “binary (RLE)”, which results in a further speedup of up to 56.3×56.3$\times$56.3 ×.
@@ 6.3.4 Scale Up the Number of Labels
Figure 10 illustrates the execution time of filtering conditions with varying numbers of labels, focusing on BL and MA, which are selected from different datasets (Neo4j and OGB) and have a relatively large number of labels. We test the filtering with $i$ labels by combining the first $i$ labels in a graph using the OR operator as the condition, to ensure that there are vertices satisfying the condition. To reduce experimental noise, we conduct the experiments for $100$ times and report the total execution time for BL and the average time for MA.
As shown in the figure, GraphAr consistently outperforms others on BL. While on MA, it performs best when the number of involved labels is no more than $40$. As the number of labels continuously increases, it performs worse than the baseline “binary (RLE)”, which is due to the number of merged intervals also increasing. In the worst case, the UDF is called for each vertex, means any two consective vertices have different labels. Considering the overhead of merging intervals from different columns, our method may perform worse than directly evaluating the UDF for each vertex. Fortunately, our investigation of real-world workloads reveals that the number of filtered labels in user-written queries is often limited. For instance, in the documentation of Neo4j examples  [14], the filtering involves no more than $5$ labels. Consequently, the performance of our method remains highly promising.
Figure 10: Execution time with the number of labels increases on, tested on BL (left) and MA (right).
@@ 6.4 Storage Media
We assess the efficiency of GraphAr across various storage media: local in-memory tmpfs, ESSD (an Alibaba Cloud virtualized elastic block device), and S3-like Object Store Service (OSS). The graph used is SF100, with specifically foucus on the comment vertex type and comment_hasTag_tag edge type.
Table 2 encapsulates the efficiency of GraphAr across different storage media. These results conclusively establish that GraphAr is not only efficient but also robust, delivering consistently high performance, with speedups of $\times8$ to 154×154$\times$154 × for neighbor retrieval and 2.7×2.7$\times$2.7 × to $\times1$ for label filtering.
6.5 End-to-end Workloads
To demonstrate the practicality of GraphAr in real-world scenarios, we conduct a performance evaluation using end-to-end workloads from the LDBC SNB benchmark [32, 57], which is widely used in the graph processing community. Although the benchmark specifies vertex/edge types, it does not explicitly define the labels. However, we are able to identify certain vertex types that are static (e.g., tagclass and place), which have a fixed and very small vertex set size that does not scale with the graph size. On the other hand, vertex types like comment and person are considered dynamic. Based on this observation, we can treat information related to static types as labels for dynamic types in GraphAr, for example, all tag classes of a comment are attached as labels for the corresponding comment vertex. Similar strategies are also adopted by graph databases [37] to optimize data access performance.
For evaluation, graphs at different scales (listed in Table 1) are generated using the LDBC SNB data generator, as Parquet files. These graphs are then converted into GraphAr format, with the vertex labels attached as described above. Upon investigating the benchmark, including $7$ short and $14$ complex interactive queries, as well as $20$ business intelligence queries, we find that neighbor retrieval is frequently encountered, involved in approximately $90\%$ of the queries. Considering the aforementioned label organization, label filtering is also common, involved in approximately $50\%$ of the queries.
@@ 6.5.1 Query Implementations
Queries. The evaluation focuses on three representative queries, with the required parameters set according to the reference implementations [12, 11]. The first query, IS-3 (interactive-short-3), aims to find all the friends of a given person and return their information. It exemplifies the common pattern of querying neighboring vertices and retrieving associated properties. The second query, IC-8 (interactive-complex-8), is more complex as it involves traversing multiple hops from the starting vertex. Lastly, the BI-2 (business-intelligence-2) query aims to analyze tag evolution. It involves finding and counting the messages associated with tags within a specific tag class, thus requiring vertex filtering by labels.
GraphAr. We develop hand-written implementations for each query based on GraphAr, which utilize the data organization and specifically prioritize two essential operations: neighbor retrieval and label filtering. Our implementation refers to the official reference implementations [12, 11] in Cypher or SQL, to ensure equivalence to the original queries.
Acero. We also implement these queries in Acero [1], which is a powerful C++ library integrated into Apache Arrow for analyzing large streams of data. It offers a comprehensive set of operators such as scan, filter, project, aggregate, and join, among others. Moreover, Acero supports taking Parquet as the data source and enables the pushdown of predicates, making it a strong baseline for comparison with GraphAr. Despite our best efforts to optimize it, we do not perform data re-organization or utilize GraphAr’s encoding/decoding optimizations for the implementations based on Acero.
@@ 6.5.2 Performance Comparison
Table 3 presents a comparison of end-to-end time performance, demonstrating that the implementation based on GraphAr significantly outperforms Acero, achieving an average speedup of 29.5×29.5$\times$29.5 ×. Upon closer examination of the results, it becomes apparent that the performance improvement can be attributed to the following factors: 1) data layout design and encoding/decoding optimizations we proposed, to enable efficient neighbor retrieval (IS-3, IC-8, BI-2) and label filtering (BI-2), as demonstrated in micro-benchmarks; 2) bitmap generation during the two critical operations, which can be utilized in subsequent selection steps (IS-3, IC-8, BI-2).
Taking IS-3 as an example, the optimization of neighbor retrieval achieves an average speedup of 571.4×571.4$\times$571.4 × compared to the baseline utilizing Parquet’s filter pushdown. The subsequent selection based on bitmaps provides a speedup of 2.7×2.7$\times$2.7 ×. Other steps like order_by show similar performance in GraphAr and Acero due to closely matched implementations.
Table 3: Execution time for IS-3, IC-8 and BI-2 (in seconds).
@@
Summary. In the comprehensive evaluation, GraphAr has been demonstrated to be a highly effective storage scheme for LPGs in data lakes. The key takeaways are:
  • Storage efficiency: GraphAr remarkably reduces storage requirements, using only $27.3\%$ of the storage compared to baseline methods for storing topology, and as low as $2.9\%$ for label storage on average.
  • Query performance: GraphAr significantly outpaces the baselines in retrieval time, achieving speedups ranging from 49.8×49.8$\times$49.8 × to 9232.4×9232.4$\times$9232.4 × for neighbor retrieval and an average speedup of 6.0×6.0$\times$6.0 × for simple label filtering, as observed in micro-benchmarks on the ESSD storage.
@@
  • Real-world relevance: In end-to-end workloads using the LDBC SNB benchmark, GraphAr shows an average speedup of 29.5×29.5$\times$29.5 × over the Acero baseline, substantiating its practical utility in real-world scenarios.